OpenStudy (anonymous):

Yellow Taxi Cabs charges a $7 flat rate in addition to $2.25 per mile. Kate has $100 to spend on transportation. How many miles can Kate travel without exceeding her limit?

OpenStudy (anonymous):

@notamathgenius

OpenStudy (anonymous):

hard

OpenStudy (anonymous):

can yall help me ?

OpenStudy (anonymous):

So it sounds like he charges 7 dollars no matter the amount of miles you go right?

OpenStudy (anonymous):

so would i just divide 100 by 7 ?

OpenStudy (anonymous):

Noooo no no no

OpenStudy (anonymous):

Do you know anything about functions?

OpenStudy (anonymous):

not really. Im taking an online class and i have no teacher

OpenStudy (anonymous):

Ok. Well you understand he charges 2.25 per mile right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

And he charges 7 dollars just for riding in his cab

OpenStudy (anonymous):

yes i understand

OpenStudy (anonymous):

Ok so x in this equation I am about to type is the number of miles

OpenStudy (anonymous):

\[2.25x+7\]

OpenStudy (anonymous):

so where x is im going to put the 100 ?

OpenStudy (anonymous):

Not quite

OpenStudy (anonymous):

will i set the equation to 0 and then solve for x ?

OpenStudy (anonymous):

\[100 = 2.25x + 7\]

OpenStudy (anonymous):

x = 41.333333333 ?

OpenStudy (anonymous):

\[100=2.25x+7\] \[100 -7=2.25x+7-7\] \[93=2.25x\] \[(93)/2.25=(2.25x)/2.25\] \[41.33=x\]

OpenStudy (anonymous):

thank you sooo much !

OpenStudy (anonymous):

can you see if you can help me on some other math problems ?

OpenStudy (anonymous):

you divide 100 by 2.25 and you will get......

OpenStudy (anonymous):

why would you divide 100 by 2.25

OpenStudy (anonymous):

shure I can help

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I just posted it on my wall

OpenStudy (anonymous):

wat wall

OpenStudy (anonymous):

my page